A rock falls from a tower that is 53.9 m high. As it is falling, its height is given by the formula h = 53.9 - 4.9t^2. How many seconds will it take for the rock to hit he ground (h=0)? Round to the nearest tenth, if necessary.

A) 7 sec

B) 592.9 sec

C) 7.3 sec

D) 3.3 sec

Answer by checkley71(8403)

0=53.9-4.9t~2 or 0=11-t2 or t~2=11 or t=3.3 sec